OpenStudy (anonymous):

please help 4tan(2arctan1/3) solve

terenzreignz (terenzreignz):

Two things to keep in mind.... First: \[\Large \tan(2\theta) = \frac{2\tan(\theta)}{1- \tan^2(\theta)}\] Second: \[\Large \tan(\arctan x )=x\]

OpenStudy (anonymous):

then?

terenzreignz (terenzreignz):

Then? Expand \(\large \tan(2\arctan \frac13)\) using \[\Large \tan(2\theta) = \frac{2\tan(\theta)}{1- \tan^2(\theta)}\] And taking \(\large \theta = \arctan\frac13\)